OpenStudy (anonymous):

Which of the following circles have their centers in the second quadrant? Check all that apply. A. (x - 5)2 + (y - 6)2 = 25 B. (x + 1)2 + (y - 7)2 = 16 C. (x - 4)2 + (y + 3)2 = 32 D. (x + 2)2 + (y - 5)2 = 9 PLEASE HELP AND EXPLAIN!!!

OpenStudy (anonymous):

The 2nd quad is (-,+) so

OpenStudy (anonymous):

Take (x-h) and (y-k) and set them both equal to 0 and solve for x and y

OpenStudy (anonymous):

For example, a: x - 5 = 0 x = 5 y - 6 = 0 y = 6 (5,6) since they are both positive this is in the first quad. So A is not right Make sense?

OpenStudy (anonymous):

Yeah thanks bro!!

OpenStudy (anonymous):

No problem, what did you get for the answer?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Well C would give you (+,-) you need (-,+)